OpenStudy (anonymous):

is (-4,7) the intersect of both lines when solving for x + 2y = 10, and 3x + 4y = 8?

terenzreignz (terenzreignz):

This is solving systems of linear equations. x + 2y = 10 3x + 4y = 8 Any idea how to go about this?

OpenStudy (anonymous):

Yeah x + 2y = 10 the answer is y = 5 (0,5) and x = 10 (10,0) for 3x + 4y = 8 the answer is (0,2) for y and the slope is -3/4. I'm just wondering if (-4, 7) is the correct intersect on both lines

terenzreignz (terenzreignz):

Oh... Graphing is a terrible way to solve systems, it's too... prone to human error. Here's an easier way :) Substitute -4 for x, and 7 for y. In the first equation x + 2y = 10 substituting -4 + 2(7) = 10 Is it valid?

OpenStudy (anonymous):

-4 + 2 * 7 = 10 yes its true

terenzreignz (terenzreignz):

Okay, so what does that tell you? It tells you that the point (-4 , 7) is on the line x + 2y = 10 But is it also on the line 3x + 4y = 8 ? Substitute and see.

OpenStudy (anonymous):

3 * -4 + 4 * 7 = 8 false, the answer is 16

terenzreignz (terenzreignz):

Then that can only mean that the point (-4, 7) is not on the line 3x + 4y = 8 What can you conclude now?

OpenStudy (anonymous):

(-4,7) isn't going the intersect of both equations

terenzreignz (terenzreignz):

Precisely :)

OpenStudy (anonymous):

How do I find the intersect of both equations then?

terenzreignz (terenzreignz):

Oh, well, let's go back to basics :D Remember your properties of equality?

OpenStudy (anonymous):

nope

terenzreignz (terenzreignz):

Okay, let a, b, c, d be real numbers If a = b then ca = cb correct?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

also, if a = b and c = d Then a + c = b + d Right? We just added the left sides and the right sides of the equations

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

have you done metrics for 2x2 system it is a simple as solving by substitution

OpenStudy (anonymous):

as simple as*

terenzreignz (terenzreignz):

x + 2y = 10 3x + 4y = 8 We can add the left sides and the right sides, and it gives us x + 3x + 2y + 4y = 18 4x + 6y = 18 But that's not really that helpful, is it? :P

OpenStudy (chihiroasleaf):

have you learn substitution or elimination method?

terenzreignz (terenzreignz):

This is the elimination method, by the way.

OpenStudy (anonymous):

What do we do after we add both sides?

terenzreignz (terenzreignz):

We don't. I'm just demonstrating adding both sides right away, and how useless it is :) Now remember, that first property I showed you, where we multiply the same number on both sides of the equation?

OpenStudy (anonymous):

Do you know what the problem I'm doing is called so I can find a video on www.khanacademy.org and see live examples?

terenzreignz (terenzreignz):

Systems of Linear Equations.